OpenStudy (anonymous):

The volume of a cylindrical can is 15.7 in.3. The diameter is 2 in. Find the height.

OpenStudy (anonymous):

radius = 1 in v=pi*r*r*h 15.7=3.142*1*1*h 15.7/(3.142*1*1)=h

OpenStudy (yttrium):

Yes. You got it right! Don't have to post it here. Hahaha!